OpenStudy (anonymous):

A sail is in the form of a right triangle. The angle at its lower edge is 57o. Which is the size of the angle formed at the top of the sail?

OpenStudy (anonymous):

is the illustration like this? 

OpenStudy (anonymous):

if it is: first a triangle has the sum of 180o angle (meaning angle1 + angle2 + angle3 = 180) second a RIGHT TRIANGLE has a 90o angle therefore: 180 = 90 + 57 + x (where x is the 3rd angle) x = 180 - 90 - 57 x = 33 [is it right?]